In the production of Zinc sulfide, 36.8 g of zinc is made to react with 19.4 g of sulfur. 

Zn + S -> ZnS

a. How many grams of ZnS is produced when zinc is completely used up? answer: 54.57 g Zns

b. Which reactant is the limiting reagent? answer: Zn

c. How many grams of the excess reagent are left? answer: 1.28 g S


solution:


1
Expert's answer
2021-12-02T01:00:31-0500

Molar mass of ZnS=97.474


Molar mass of Zn=65.38

36.8/65.38=0.56

Molar mass of S= 32.065

19.4/31.065=0.624

= (0.624+0.56)97.472=115.46g